# Holistic Order Prediction in Natural Scenes

### v1

@@ C.1 Benchmark specifications
We specify the backbones used for our main benchmark (table˜1) of the main paper in table˜3. Note that Mask2Order and InstaFormer’s backbones are frozen at training time.
@@ C.2 LLaVA
Benchmark. For our benchmark (table˜1, in the main paper), we use LLaVA 1.5 Liu et al. (2024a), which is composed of a CLIP ViT-L 336 pixels vision encoder and a Vicuna 7B Chiang et al. (2023) token generator.
@@
In both cases, we prompt the model with the InstaOrder-Vqa evaluation set and report the results using the same metrics as with the other methods. A description of our pipeline is detailed in fig.˜6 of the supplementary material.
@@ C.3 Foundation models
We use Grounded SAM Ren et al. (2024); Kirillov et al. (2023) and Depth Anything V2 Yang et al. (2024b) to obtain instance-level depth estimations that we can then compare together to obtain depth ordering predictions. A diagram of the overall pipeline is depicted in fig.˜7. To propose a baseline as competitive as possible, we select the best-performing networks available at the time of writing, i.e., Grounding DINO Swin-T [^1] The Grounded DINO model for Grounded SAM is available at this URL., SAM ViT-H [^2] The SAM model zoo is available at this URL. and Depth Anything V2 ViT-L [^3] The Depth Anything V2 model zoo is available at this URL..
@@
Depth order heuristics. There are several ways to obtain instance-wise depth orderings given segment-level depth maps. In our work and benchmark, we report three heuristics that yield different results (reported in table˜1, in the main paper): min-max, mean, and median.
@@
Subsequently, sorting the resulting ranges provides the desired depth orderings. In practice, the min-max heuristic is slightly flawed as the instance masks are not always perfectly aligned with the instance-level depth maps. This can result in obtaining minimum or maximum values slightly outside the actual instances, generating bias in the sorting process. Instead, we find the other two heuristics simpler and more robust to noise (as results reveal in table˜1 of the main paper).
@@
In practice, we find the median heuristic to perform the best (cf. table˜1, in the main paper).
Appendix D Evaluation metrics
D.1 Occlusion order
Following previous works Zhan et al. (2020); Zhu et al. (2017); Lee and Park (2022), we evaluate the performance of our networks on the task of occlusion ordering prediction by reporting precision, recall, and F1-score. The predictions are evaluated by comparing each element of the predicted ordering matrix to each element of the ground truth matrix. Given two instances $A$ and $B$, the aforementioned metrics are expressed as follows:
D.2 Depth order
where $w_{AB}=\frac{2}{\text{count}_{AB}}$.
Appendix E Adapters
@@
We conduct experiments to study the effects of these components on InstaFormer in section˜F.2.
@@
Appendix F Ablations studies
F.1 Segmentation backbone size
@@
Occlusion order recovery. We report the results when varying the backbone of the segmentation network from Swin-T100 to Swin-L200 in table˜4(a). For Mask2Ordero, we observe a noticeable performance increase when using IN-21K pre-trained backbones. We observe a similar trend for Mask2Ordero,d.
@@
Depth order recovery. We report the results when varying the backbone of the segmentation backbone from Swin-T100 to Swin-L200 in table˜4(b). These results show very different trends with respect to the occlusion order backbone ablation. For Mask2Ordero networks, they seem to hastily plateau around 18.75 WHDR (all) with slight improvement from Swin-T100 to Swin-B100. Increasing the number of queries from 100 to 200 slightly fosters performance. We note the positive effect of simultaneously performing both the task of occlusion and depth predictions as the results of Mask2Ordero,d Swin-T100 already overcomes those of Mask2Ordero Swin-L200. Moreover, Mask2Ordero,d Swin-L200 beats its Mask2Ordero counterpart by 1.43 points of WHDR (all).
Task difficulty. We observe that the holistic task of depth order prediction seems to be easier to perform than the task of occlusion order prediction since the WHDR discrepancy from InstaOrderNets and InstaFormer is larger than their F1 difference (cf. table˜1 of the main paper) and since changing from a weaker to a stronger backbone mostly benefits the occlusion order prediction networks (from table˜4(a) and table˜4(b)).
@@
F.2 Adapters
We ablate the adapters of InstaFormer models. The results of these experiments are reported in table˜5.
@@ F.2 Adapters
Appendix G Inference cost
G.1 Runtime comparison
We benchmark the runtime inference of our holistic InstaFormer and compare it against pairwise InstaOrderNet, in fig.˜5(a) of the main paper. Runtime for our proposed InstaFormer appears constant with respect to the number of instances in the image, while InstaOrderNet is exponential. Moreover, InstaFormer’s efficiency is amortized for any image containing more than 7 instances and up to 12 times faster than the baseline when the image contains 20 instances. This makes our approach suitable for prediction on natural images since such scenes usually contain many objects.
@@ G.1 Runtime comparison
G.2 Memory consumption
We further benchmark the memory consumption of our InstaFormero,d network and compare it to InstaFormero,d in fig.˜5(b) of the main paper. The memory consumption of our InstaFormer grows linearly with the number of instances on the image, while InstaOrderNet’s remains constant. This highlights the trade-off between inference speed and memory consumption. We argue that InstaFormer still wins that trade-off since its runtime is constant while InstaOrderNet’s is exponential.
@@ G.2 Memory consumption
G.3 Scalability
In table˜6, we compute the number of average instances detected in images as well as the number of maximum predicted instances in a single sample across the validation set of COCO Lin et al. (2014). We vary the number of object queries, but we observe in both cases that the number of instances predicted in a sample is close or above to 11. Comparing this result to fig.˜5(a) and fig.˜5(b), we conclude that it is on average better to use InstaFormer than other baselines since natural images contain, on average, close to, or above 11 instances.
The second column of table table˜6 emphasizes the fact that our InstaFormer network can scale way above 10 instances although the training set of InstaOrder is limited to annotations containing 10 objects Lee and Park (2022). As a matter of fact, InstaFormer with 200 object queries detects 63 instances on an image of the COCO validation set. In both 100 and 200 object queries settings, the value of maximum instances detected is far below the total number of object queries. This leads us to conclude that InstaFormer is practical for general inferences in natural scenes.
@@ G.3 Scalability
Appendix H Qualitative results
H.1 Comments on results from the main paper
@@
In fig.˜1 of the main paper, we display two in-the-wild results. In the “athletics” example, we note the occlusion from the hand of “Person1” to “Person3” that was effectively captured from our model. Note that this occlusion is difficult to notice without carefully observing the input RGB image.
@@
Analyzing fig.˜4, in the main paper, the first row contains “elephants” and “K-Pop”, which are both in-the-wild predictions from our network. We observe an ambiguous occlusion case in the “elephants” example, where “Elephant3” does not seem to occlude “Elephant2” even though reported as such on the occlusion graph. While it is ambiguous whether “Elephant3” is closer to the camera than “Elephant1”, the network still appears to predict a coherent order.
The second image of the first row of fig.˜4, namely “K-Pop” is an impressive example of accurate depth order prediction in a complex scene. While the input RGB image solely depicts silhouettes, our InstaFormer network still predicts an accurate and coherent depth layout for the scene. In that case, the network shows difficulties understanding the occlusion orderings for some elements since there are little clues to rely on. For example, “Person6” and “Person4” are linked with a bidirectional occlusion even though it seems that it should be unidirectional from “Person4” to “Person6”. The same holds for “Person1” and “Person3”.
Figure 8: Ground-truth visualization of the predictions from fig.˜4. We visualize the order graphs and the masks from the InstaOrder dataset and overlap their respective COCO labels on top of them. However, we re-use the same colors as the predicted classes from InstaFormer in fig.˜4 for better comparison with the predictions. We use a white box with colored border to indicate ground-truth that have not been matched from our network’s predictions. Annotated objects on the image not present in the graphs symbolizes that our network predicted these nodes even though not present in the ground truth.
All the remaining images of fig.˜4 are extracted from the validation set of InstaOrder Lee and Park (2022). We specifically highlight the “donuts” example to emphasize that the model can still predict accurate orderings even crowded scenes.
@@
On the last row of fig.˜4, the “cooking cat” example shows that our model can predict sane relations even if the “Cat” is blurred. As a matter of fact, the cat indeed occludes the “Oven”, yet the “Oven” does not. Careful inspection reveals that the blurriness of the cat lets appear a notch of the “Oven” that visually occludes the “Cat” even though it is physically impossible.
Finally, the last example on the last row and last column of fig.˜4, called “stop sign”, shows a neat case of a prediction of perfect occlusion and depth orders, even though the size of the cars is noticeably small. We also emphasize this result as no occlusion between elements occurs. We recall that in COCO Lin et al. (2014), the stop sign class is only annotated on the sign itself, and not the pole.
H.2 Ground-truth visualizations
We visualize the ground-truths for the InstaOrder validation set prediction of fig.˜4 in fig.˜8 and fig.˜10 on fig.˜11. Note that, in some rare cases, our network was unable to detect some of the instances in the image, leading to no ordering prediction (fig.˜8, “donuts", “cat in the oven"). However, more frequently, it predicted more instances than the ground-truth annotation, for example, “sheep2” is accurately detected along with its proper occlusion ordering (first row, last column of fig.˜10 and fig.˜11). This is also the case for the teddy bears (see fig.˜10 and fig.˜11). Sometimes, our network predicts even better orderings than the ground-truth, for example, in fig.˜10, in the cake picture (third row, second column), the network predicts a bidirectional occlusion between “Person3" and “Cake". This seems more reasonable than the ground-truth annotating a unidirectional order (fig.˜11). In some cases, the objects are too thin to be detected, in which case the predicted orderings are still sane, although they could probably be more precise (“kite", 4th row, 2nd column on fig.˜10 and fig.˜11)
(a) Instance-mixup.
(b) Segmentation failure.
Figure 9: Failure cases of InstaFormero,d. We highlight two failure cases of our network: the “instance-mixup” and the “segmentation failure”. In the “instance-mixup” case, the “Person” instances related to the “Ski” are shared instead of being targeted solely on “Person1”. In the “segmentation failure” case, a singular instance “Elephant” is wrongfully segmented multiple times, leading to multiple false positive nodes in the geometrical graphs.
H.3 Failure case
We highlight two cases in which InstaFormer struggles to predict accurate relationships between instances. We name these situation “instance mixup” (fig.˜9(a)) and “segmentation failure” (fig.˜9(b)).
H.4 More qualitative results
We provide even more qualitative results obtained from InstaFormero,d on InstaOrder in fig.˜10. And report their ground-truths in fig.˜11.
@@ H.4 More qualitative results
Figure 11: Ground-truth visualization of the predictions from fig.˜10. We visualize the order graphs and the masks from the InstaOrder dataset and overlap their respective COCO labels on top of them. However, we re-use the same colors as the predicted classes from InstaFormer in fig.˜10 for better comparison with the predictions. We use a white box with colored border to indicate ground-truth that have not been matched from our network’s predictions.
Appendix I Web image sources

### v2

@@ C.1 Benchmark specifications
We specify the backbones used for our main benchmark (Tab. 1) of the main paper in Tab. 3. Note that Mask2Order and InstaFormer’s backbones are frozen at training time.
@@ C.2 LLaVA
Benchmark. For our benchmark (Tab. 1, in the main paper), we use LLaVA 1.5 Liu et al. (2024a), which is composed of a CLIP ViT-L 336 pixels vision encoder and a Vicuna 7B Chiang et al. (2023) token generator.
@@
In both cases, we prompt the model with the InstaOrder-Vqa evaluation set and report the results using the same metrics as with the other methods. A description of our pipeline is detailed in Fig. 7 of the supplementary material.
@@ C.3 Foundation models
We use Grounded SAM Ren et al. (2024); Kirillov et al. (2023) and Depth Anything V2 Yang et al. (2024b) to obtain instance-level depth estimations that we can then compare together to obtain depth ordering predictions. A diagram of the overall pipeline is depicted in Fig. 8. To propose a baseline as competitive as possible, we select the best-performing networks available at the time of writing, i.e., Grounding DINO Swin-T [^1] The Grounded DINO model for Grounded SAM is available at this URL., SAM ViT-H [^2] The SAM model zoo is available at this URL. and Depth Anything V2 ViT-L [^3] The Depth Anything V2 model zoo is available at this URL..
@@
Depth order heuristics. There are several ways to obtain instance-wise depth orderings given segment-level depth maps. In our work and benchmark, we report three heuristics that yield different results (reported in Tab. 1, in the main paper): min-max, mean, and median.
@@
Subsequently, sorting the resulting ranges provides the desired depth orderings. In practice, the min-max heuristic is slightly flawed as the instance masks are not always perfectly aligned with the instance-level depth maps. This can result in obtaining minimum or maximum values slightly outside the actual instances, generating bias in the sorting process. Instead, we find the other two heuristics simpler and more robust to noise (as results reveal in Tab. 1 of the main paper).
@@
In practice, we find the median heuristic to perform the best (cf. Tab. 1, in the main paper).
Appendix D Adapters
@@
We conduct experiments to study the effects of these components on InstaFormer in Section˜E.2.
@@
Appendix E Ablations studies
E.1 Segmentation backbone size
@@
Occlusion order recovery. We report the results when varying the backbone of the segmentation network from Swin-T100 to Swin-L200 in Tab. 4(a). For Mask2Ordero, we observe a noticeable performance increase when using IN-21K pre-trained backbones. We observe a similar trend for Mask2Ordero,d.
@@
Depth order recovery. We report the results when varying the backbone of the segmentation backbone from Swin-T100 to Swin-L200 in Tab. 4(b). These results show very different trends with respect to the occlusion order backbone ablation. For Mask2Ordero networks, they seem to hastily plateau around 18.75 WHDR (all) with slight improvement from Swin-T100 to Swin-B100. Increasing the number of queries from 100 to 200 slightly fosters performance. We note the positive effect of simultaneously performing both the task of occlusion and depth predictions as the results of Mask2Ordero,d Swin-T100 already overcomes those of Mask2Ordero Swin-L200. Moreover, Mask2Ordero,d Swin-L200 beats its Mask2Ordero counterpart by 1.43 points of WHDR (all).
Task difficulty. We observe that the holistic task of depth order prediction seems to be easier to perform than the task of occlusion order prediction since the WHDR discrepancy from InstaOrderNets and InstaFormer is larger than their F1 difference (cf. Tab. 1 of the main paper) and since changing from a weaker to a stronger backbone mostly benefits the occlusion order prediction networks (from Tab. 4(a) and Tab. 4(b)).
@@
E.2 Adapters
We ablate the adapters of InstaFormer models. The results of these experiments are reported in Tab. 5.
@@ E.2 Adapters
Appendix F Inference cost
F.1 Runtime comparison
We benchmark the runtime inference of our holistic InstaFormer and compare it against pairwise InstaOrderNet, in Fig. 5(a) of the main paper. Runtime for our proposed InstaFormer appears constant with respect to the number of instances in the image, while InstaOrderNet is exponential. Moreover, InstaFormer’s efficiency is amortized for any image containing more than 7 instances and up to 12 times faster than the baseline when the image contains 20 instances. This makes our approach suitable for prediction on natural images since such scenes usually contain many objects.
@@ F.1 Runtime comparison
F.2 Memory consumption
We further benchmark the memory consumption of our InstaFormero,d network and compare it to InstaFormero,d in Fig. 5(b) of the main paper. The memory consumption of our InstaFormer grows linearly with the number of instances on the image, while InstaOrderNet’s remains constant. This highlights the trade-off between inference speed and memory consumption. We argue that InstaFormer still wins that trade-off since its runtime is constant while InstaOrderNet’s is exponential.
@@ F.2 Memory consumption
F.3 Scalability
In Tab. 6, we compute the number of average instances detected in images as well as the number of maximum predicted instances in a single sample across the validation set of COCO Lin et al. (2014). We vary the number of object queries, but we observe in both cases that the number of instances predicted in a sample is close or above to 11. Comparing this result to Fig. 5(a) and Fig. 5(b), we conclude that it is on average better to use InstaFormer than other baselines since natural images contain, on average, close to, or above 11 instances.
The second column of Tab. 6 emphasizes the fact that our InstaFormer network can scale way above 10 instances although the training set of InstaOrder is limited to annotations containing 10 objects Lee and Park (2022). As a matter of fact, InstaFormer with 200 object queries detects 63 instances on an image of the COCO validation set. In both 100 and 200 object queries settings, the value of maximum instances detected is far below the total number of object queries. This leads us to conclude that InstaFormer is practical for general inferences in natural scenes.
@@ F.3 Scalability
Appendix G Qualitative results
G.1 Comments on results from the main paper
@@
In Fig. 1 of the main paper, we display two in-the-wild results. In the “athletics” example, we note the occlusion from the hand of “Person1” to “Person3” that was effectively captured from our model. Note that this occlusion is difficult to notice without carefully observing the input RGB image.
@@
Analyzing Fig. 4, in the main paper, the first row contains “elephants” and “K-Pop”, which are both in-the-wild predictions from our network. We observe an ambiguous occlusion case in the “elephants” example, where “Elephant3” does not seem to occlude “Elephant2” even though reported as such on the occlusion graph. While it is ambiguous whether “Elephant3” is closer to the camera than “Elephant1”, the network still appears to predict a coherent order.
The second image of the first row of Fig. 4, namely “K-Pop” is an impressive example of accurate depth order prediction in a complex scene. While the input RGB image solely depicts silhouettes, our InstaFormer network still predicts an accurate and coherent depth layout for the scene. In that case, the network shows difficulties understanding the occlusion orderings for some elements since there are little clues to rely on. For example, “Person6” and “Person4” are linked with a bidirectional occlusion even though it seems that it should be unidirectional from “Person4” to “Person6”. The same holds for “Person1” and “Person3”.
Figure 9: Ground-truth visualization of the predictions from Fig. 4. We visualize the order graphs and the masks from the InstaOrder dataset and overlap their respective COCO labels on top of them. However, we re-use the same colors as the predicted classes from InstaFormer in Fig. 4 for better comparison with the predictions. We use a white box with colored border to indicate ground-truth that have not been matched from our network’s predictions. Annotated objects on the image not present in the graphs symbolizes that our network predicted these nodes even though not present in the ground truth.
All the remaining images of Fig. 4 are extracted from the validation set of InstaOrder Lee and Park (2022). We specifically highlight the “donuts” example to emphasize that the model can still predict accurate orderings even crowded scenes.
@@
On the last row of Fig. 4, the “cooking cat” example shows that our model can predict sane relations even if the “Cat” is blurred. As a matter of fact, the cat indeed occludes the “Oven”, yet the “Oven” does not. Careful inspection reveals that the blurriness of the cat lets appear a notch of the “Oven” that visually occludes the “Cat” even though it is physically impossible.
Finally, the last example on the last row and last column of Fig. 4, called “stop sign”, shows a neat case of a prediction of perfect occlusion and depth orders, even though the size of the cars is noticeably small. We also emphasize this result as no occlusion between elements occurs. We recall that in COCO Lin et al. (2014), the stop sign class is only annotated on the sign itself, and not the pole.
G.2 Ground-truth visualizations
We visualize the ground-truths for the InstaOrder validation set prediction of Fig. 4 in Fig. 9 and Fig. 10 on Fig. 11. Note that, in some rare cases, our network was unable to detect some of the instances in the image, leading to no ordering prediction (Fig. 9, “donuts", “cat in the oven"). However, more frequently, it predicted more instances than the ground-truth annotation, for example, “sheep2” is accurately detected along with its proper occlusion ordering (first row, last column of Fig. 10 and Fig. 11). This is also the case for the teddy bears (see Fig. 10 and Fig. 11). Sometimes, our network predicts even better orderings than the ground-truth, for example, in Fig. 10, in the cake picture (third row, second column), the network predicts a bidirectional occlusion between “Person3" and “Cake". This seems more reasonable than the ground-truth annotating a unidirectional order (Fig. 11). In some cases, the objects are too thin to be detected, in which case the predicted orderings are still sane, although they could probably be more precise (“kite", 4th row, 2nd column on Fig. 10 and Fig. 11)
G.3 More qualitative results
We provide even more qualitative results obtained from InstaFormero,d on InstaOrder in Fig. 10. And report their ground-truths in Fig. 11.
@@ G.3 More qualitative results
Figure 11: Ground-truth visualization of the predictions from Fig. 10. We visualize the order graphs and the masks from the InstaOrder dataset and overlap their respective COCO labels on top of them. However, we re-use the same colors as the predicted classes from InstaFormer in Fig. 10 for better comparison with the predictions. We use a white box with colored border to indicate ground-truth that have not been matched from our network’s predictions.
Appendix H Web image sources